OpenStudy (anonymous):

Mrs. AYer IS PAINTING THE OUTSIDE OF HER SON'S TOY BOX, INCLUDING THE TOP AND BOTTOM. tHE tOY BOX MEASURES 3 FEET LONG, 1.5 FEET WIDE, AND 2 FEET HIGH. whAT IS THE TOTAL SURFACE AREA SHE WILL PAINT? A) 9.0 FT^2 B) 13.5 FT^2 C) 22.5 FT^2 D) 27.0 FT^2

OpenStudy (anonymous):

SO WOULD YOU HAVE TO MULTIPLE THE DIFFERENT MEASUREMENTS?

OpenStudy (anonymous):

3*1.5=4.5ft 2*4.5 = 9ft ( Area bottom an top ) 3*2 = 6 2*6=12 (Area bottom 2 side) 1.5*2=3ft 2 * 3 = 6 (Area front an back) 9+12+6=27ft

OpenStudy (anonymous):

wait I was wrong

OpenStudy (anonymous):

3*1.5=4.5ft 2*4.5 = 9ft ( Area bottom an top ) 3*2 = 6 2*6=12 (Area 2 side) 1.5*2=3ft 2 * 3 = 6 (Area front an back) 9+12+6=27ft

OpenStudy (anonymous):

D) If I deserve medal then give me one

OpenStudy (anonymous):

u were right:)